OpenStudy (anonymous):

brandy is making a peach pie for her grandmother for Christmas. The diameter of the pie is 12 in. If three-fourths of the area of the pie was eaten after Christmas dinner, what is the area of the top of the pie that is left over? A.113.10in^2 B.56.55in^2 C.28.27in^2 D.226.19in^2

OpenStudy (anonymous):

three fourths were eaten, then we have still one forth. so the area= (1/4)*r^2*pi we have r=diameter/2=6 area= (1/4)*36*pi=9 pi = c)28.26

OpenStudy (anonymous):

Thankyou!

OpenStudy (anonymous):

you are welcome mate ;)